OpenStudy (anonymous):

Find (f+g)(x) and (f-g) (x) for the given functions f and g. f(x) = x +8, g(x) = x-8

OpenStudy (anonymous):

I am trying to figure out this how set this up and solve

OpenStudy (owlcoffee):

Whenever we operate functions, we treat them as numbers, to later on, as a univariable equality. I'll do the sum, to show you an example: if: \[f:f(x)=x+8\] \[g:g(x)=x-8\] Then, the sum would be: \[(f+g)(x)=f(x)+g(x)\] but we already know the values of them, they were given there above, so let's replace it: \[(f+g)(x)=(x+8)+(x-8)\] Doing a little algebra, and getting rid of the parenthesis: \[(f+g)(x)=x+8+x-8\] \[(f+g)(x)=2x\]

OpenStudy (owlcoffee):

It's the same with the sustraction, try it and let me know the result you got :)

OpenStudy (anonymous):

Is the other one 8x

OpenStudy (anonymous):

@Owlcoffee is this right

OpenStudy (owlcoffee):

That's, incorrect, friend, can I see what you did?

OpenStudy (anonymous):

I deleted what I did already

OpenStudy (owlcoffee):

Well, don't worry, I'll give you a little step: if: \[f:f(x)=x+8\] \[g:g(x)=x-8\] Then: \[(f-g)(x)=(x+8)-(x-8)\] remember the distributive property, or when you sustracted polynomials earlier in basic algebra, so all we do is distribute that "-" and by the law of signs: \[(f-g)(x)=x+8-x+8\] I'll let you handle the rest.

OpenStudy (anonymous):

is 16 the answer

OpenStudy (owlcoffee):

that's correct.

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

Thanks for your help

OpenStudy (anonymous):

Will it be 16x or just 16 @Owlcoffee

OpenStudy (owlcoffee):

just 16, because the x's cancelled. we call that a "constant function"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks for your help today